The table below shows data regarding the world indoor records in the men’s 400-m run.
YEAR RECORDS IN THE 400-M RUN IN SECONDS
_________________________________________________
1970 46.8
2004 44.63
a. Use the two data points to find a linear function that fits the data. Let the number of years since 1970 and the world record x years from 1970. Show your work and explain your answer.
b. Use the function to estimate the world record in the men’s 400-m run in 2008 and in 2010.

Answer by stanbon(57967)

The table below shows data regarding the world indoor records in the men’s 400-m run.
YEAR RECORDS IN THE 400-M RUN IN SECONDS
_________________________________________________
1970 46.8
2004 44.63
a. Use the two data points to find a linear function that fits the data. Let the number of years since 1970 and the world record x years from 1970. Show your work and explain your answer.
---
You have two points: (0,46.8) and (34,44.63)
slope = (44.63-46.8)/34 = -0.0638
Equation:
record = -0.0638x + 46.8
-------------------------------
b. Use the function to estimate the world record in the men’s 400-m run in 2008 and in 2010.
2008 has x = 38
2010 has x = 40
---
record(2008) = -0.0638(38) + 46.8
record(2010) = -0.0638(40) + 46.8
=================
Cheers,
Stan H.